# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v2

@@ 5 Main Results: Bandits with Preference Feedback
We return to our main problem setting in which a pair of actions, ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$, are chosen and the observed response is a noisy binary indicator of ${\bm{x}}_{t}$ yielding a higher utility than ${\bm{x}}^{\prime}_{t}$. While this type of feedback is more consistent in practice, it creates quite a challenging problem compared to the logistic problem of Section 4. The search space for action pairs ${\mathcal{X}}\times{\mathcal{X}}$ is significantly larger than ${\mathcal{X}}$, and the observed preference feedback of $s(f({\bm{x}}_{t})-f({\bm{x}}^{\prime}_{t}))$ conveys only relative information between two actions. We start by presenting a solution to estimate $f$ and obtain valid confidence sets under preference feedback. Using these confidence sets we then design the MaxMinLCB algorithm which chooses action pairs that are not only favorable, i.e., yield high utility, but are also informative for improving the utility confidence sets.
@@ 5.1 Preference-based Confidence Sets
We consider the probabilistic model of $y_{t}$ as stated in Section 3, and write the corresponding regularized negative loglikelihood loss as
@@
This loss may be optimized over different function classes and is commonly used for linear dueling bandits (e.g., Saha, 2021), and has been notably successful in reinforcement learning with human feedback (Christiano et al., 2017). We proceed to show that the preference-based loss ${\mathcal{L}}_{k}^{\mathrm{D}}$ is equivalent to ${\mathcal{L}}_{k^{\mathrm{D}}}^{\mathrm{L}}$, the standard logistic loss (3) invoked with a specific kernel function $k^{\mathrm{D}}$. This will allow us to cast the problem of inference with preference feedback as a kernelized logistic regression problem. To this end, we define the dueling kernel as
@@
The proof is left to Section C.1. The assumption on eigenfunctions in Proposition 4 is primarily made to simplify the equivalence class. In particular, the relative preference function $h$ can only capture the utility $f$ up to a bias, i.e., if a constant bias $b$ is added to all values of $f$, the corresponding $h$ function will not change. The value of $b$ may not be recovered by drawing queries from $h$, however, this will not cause issues in terms of identifying $\operatorname*{arg\,max}$ of $f$ through querying values of $h$. Therefore, we set $b=0$ by assuming that eigenfunctions of $k$ are zero-mean. This assumption automatically holds for all kernels that are translation or rotation invariant over symmetric domains, since their eigenfunctions are periodic $L_{2}({\mathcal{X}})$ basis functions, e.g., Matérn kernels and sinusoids.
Proposition 4 allows us to re-write the preference-based loss function of (6) as a logistic-type loss
@@
Assume $f\in{\mathcal{H}}_{k}$ and $\left\lVert f\right\rVert_{k}\leq B$. Choose $10<\delta<$ 1 and set $\beta^{\mathrm{D}}_{t}(\delta)$ and $\sigma^{\mathrm{D}}_{t}$ as in equations (4) and (5), with $k^{D}$ used as the kernel function. Then,
@@
Corollary 5 gives valid confidence sets for kernelized utility functions under preference feedback and immediately improves prior results on linear dueling bandits and kernelized dueling bandits with regression-type loss, to kernelized setting with logistic-type likelihood. To demonstrate this, in Section C.3 we present the kernelized extensions of MaxInP (Saha (2021), Algorithm 3), and IDS (Kirschner and Krause (2021), Algorithm 4) and prove the corresponding regret guarantees (cf. Theorems 15 and 16). Corollary 5 holds almost immediately by invoking Theorem 2 with the dueling kernel $k^{\mathrm{D}}$ and applying Proposition 4. A proof is provided in Section C.1 for completeness.
Comparison to Prior Work. A line of previous work assumes that both $f$ and the probability $s(f({\bm{x}}))$ are $B$-bounded members of ${\mathcal{H}}_{k}$. This allows them to directly estimate $s(f({\bm{x}}))$ via kernelized linear regression (Xu et al., 2020; Mehta et al., 2023b; Kirschner and Krause, 2021). The resulting confidence intervals are then around the least squares estimator, which does not align with the logistic estimator ${f_{t}}$. This model does not encode the fact that $s(f({\bm{x}}))$ only takes values in $[0,1]$ and considers a sub-Gaussian distribution for $y_{t}$, instead of the Bernoulli formulation when calculating the likelihood. Therefore, the resulting algorithms require more samples to learn an accurate reward estimate. In a concurrent work, Xu et al. (2024) also consider the loss function of Equation 6 and present likelihood-ratio confidence sets. The width of the sets at time $T$, scales with $\sqrt{T\log{\mathcal{N}}({\mathcal{H}}_{k};1/T})$ where the second term is the metric entropy of the $B$-bounded RKHS at resolution $1/T$, that is, the log-covering number of this function class, using balls of radius $1/T$. It is known that $\log{\mathcal{N}}({\mathcal{H}}_{k};1/T)\asymp\gamma_{T}$ as defined in Theorem 2. This may be easily verified using Wainwright (2019, Example 5.12) and (Vakili et al., 2021, Definition 1). Noting the definition of $\beta_{t}^{\mathrm{D}}$, we see that likelihood ratio sets of Xu et al. (2024) are wider than Corollary 5. Consequently, the presented regret guarantee in this work is looser by a factor of $T^{1/4}$ compared to our bound in Theorem 6.
@@
[…]
and $\text{and }{\bm{x}}_{t}^{\prime}$ $={\bm{x}}^{\prime}({\bm{x}}_{t}).$
@@ 5.2 Action Selection Strategy
Let ${\bm{x}}^{\prime}({\bm{x}})=\arg\min_{{\bm{x}}^{\prime}\in{\mathcal{M}}_{t}}\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})$ denote a response function. We propose MaxMinLCB in Algorithm 1 for the preference feedback bandit problem that selects ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ jointly as
@@
where the lower-confidence bound $\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})=s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))-\beta^{\mathrm{D}}_{t}\sigma^{\mathrm{D}}_{t}({\bm{x}},{\bm{x}}^{\prime})$ presents a pessimistic estimate of $h$ and ${\mathcal{M}}_{t}=\{{\bm{x}}\in{\mathcal{X}}|\,\forall{\bm{x}}^{\prime}\in{\mathcal{X}}:\,s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))+\beta^{\mathrm{D}}_{t}\sigma^{\mathrm{D}}_{t}({\bm{x}},{\bm{x}}^{\prime})\geq 0.5\}$ is the set of potentially optimal actions. The second action is chosen as ${\bm{x}}_{t}^{\prime}={\bm{x}}^{\prime}({\bm{x}}_{t})$. Equation 7 forms a zero-sum Stackelberg (Leader–Follower) game where the actions ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ are chosen sequentially (Stackelberg, 1952). First, the Leader selects ${\bm{x}}_{t}$, then the Follower selects ${\bm{x}}_{t}^{\prime}$ depending on the choice of ${\bm{x}}_{t}$. Importantly, due to the sequential nature of action selection, ${\bm{x}}_{t}$ is chosen by the Leader such that the Follower’s action selection function, ${\bm{x}}^{\prime}(\cdot)$, is accounted for in the selection of ${\bm{x}}_{t}$. Sequential optimization problems are known to be computationally NP-hard even for linear functions (Jeroslow, 1985). However, due to their importance in practical applications, there are algorithms that can efficiently approximate a solution over large domains (Sinha et al., 2017; Ghadimi and Wang, 2018; Dagréou et al., 2022; Camacho-Vallejo et al., 2023).
@@
MaxMinLCB builds on a simple insight: if the utility $f$ is known, both the Leader and the Follower will choose ${\bm{x}}^{\star}$ yielding an objective value $0.5$ for both players, and zero dueling regret. Since MaxMinLCB has no access to $f$, it leverages the confidence sets of Corollary 5 and uses a pessimistic approach by considering the LCB instead. There are two crucial properties of the Follower specific to this game. First, the Follower can not do worse than the Leader with respect to the $\mathrm{LCB}_{t}$. In any scenario, the Follower can match the Leader’s action which results in $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})=0.5$. Second, for sufficiently tight confidence sets, the Follower will not select sub-optimal actions. In this case, the Leader’s best action must be optimal as it anticipates the Follower’s response and Equation 7 recovers the optimal actions. Therefore, the objective value of the game considered in Equation 7 is always less than, or equal to the objective of the game with known utility function $f$, i.e., $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})\leq 0.5=f({\bm{x}}^{\star},{\bm{x}}^{\star})$ and the gap shrinks with the confidence sets. Overall, the Stackelberg game in Equation 7 can be considered as a lower approximation of the game played with known utility function $f$.
@@
The primary challenge for MaxMinLCB is to sample action pairs that sufficiently shrink the confidence sets for the optimal actions without accumulating too much regret. MaxMinLCB balances this exploration-exploitation trade-off naturally with its game theoretic formulation. We view the selection of ${\bm{x}}_{t}$ to be exploitative by trying to maximize the unknown utility $f({\bm{x}}_{t})$ and minimizing regret. On the other hand, ${\bm{x}}_{t}^{\prime}$ is chosen to be the most competitive opponent to ${\bm{x}}_{t}$, i.e., testing whether the condition $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})\geq 0.5$ holds. Note that $\mathrm{LCB}_{t}$ is pessimistic concerning ${\bm{x}}_{t}$ making it robust against the uncertainty in the confidence set estimation. At the same time, $\mathrm{LCB}_{t}$ is an optimistic estimate for ${\bm{x}}_{t}^{\prime}$ encouraging exploration. In our main theoretical result, we prove that under the assumptions of Corollary 5, MaxMinLCB achieves sublinear regret on the dueling bandit problem.
@@
Suppose the utility function $f$ lies in ${\mathcal{H}}_{k}$ with a norm bounded by $B$, and that kernel $k$ satisfies the assumption of Proposition 4. Let $\delta\in(0,1]$ and choose the exploration coefficient $\beta^{\mathrm{D}}_{t}(\delta)$ as in Corollary 5. Then MaxMinLCB satisfies the anytime dueling regret of
@@
The proof is left to Section C.2. The information gain $\gamma_{T}^{\mathrm{D}}$ in Theorem 6 quantifies the structural complexity of the RKHS corresponding to $k^{\mathrm{D}}$ and its dependence on $T$ is fairly understood for kernels commonly used in applications of bandit optimization. As an example, for a Matérn kernel of smoothness $\nu$ defined over a $d$-dimensional domain, $\gamma_{T}=\tilde{\mathcal{O}}(T^{d/(2\nu+d)})$ (Remark 2, Vakili et al., 2021) and the corresponding regret bound grows sublinearly with $T$.
Restricting the optimization domain to ${\mathcal{M}}_{t}\subset{\mathcal{X}}$ is common in the literature (Zoghi et al., 2014a; Saha, 2021) despite being challenging in applications with large or continuous domains. We conjecture that MaxMinLCB would enjoy similar regret guarantees without restricting the selection domain to ${\mathcal{M}}_{t}$ as done in Equation 7. This claim is supported by our experiments in Section 6.2 which are carried out without such restriction on the optimization domain.
@@
Figure 1: Regret of learning the Ackley function with logistic and preference feedback. (a) Same UCB algorithms, each using a different confidence set. LGP-UCB performs best, showcasing the power of Theorem 2. (b): Algorithms with different acquisition functions, all using our confidence sets. MaxMinLCB is more sample-efficient.
@@ 6 Experiments
Our experiments are on finding the maxima of test functions commonly used in (non-convex) optimization literature (Jamil and Yang, 2013), given only preference feedback. These functions cover challenging optimization landscapes including several local optima, plateaus, and valleys, allowing us to test the versatility of MaxMinLCB. We use the Ackley function for illustration in the main text and provide the regret plots for the remainder of the functions in Appendix E. For all experiments, we set the horizon $T=2000$ and evaluate all algorithms on a uniform mesh over the input domain of size $100$. Additionally, we conducted experiments on the Yelp restaurant review dataset to demonstrate the applicability of MaxMinLCB on real-world data and its scaling to larger domains. All experiments are run across $20$ random seeds and reported values are averaged over the seeds, together with standard error. The environments and algorithms are implemented [^2] The code is made available at github.com/lasgroup/MaxMinLCB. end-to-end in JAX (Bradbury et al., 2018).

### v3

@@ 5 Main Results: Bandits with Preference Feedback
We return to our main problem setting in which a pair of actions, ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$, are chosen and the observed response is a noisy binary indicator of ${\bm{x}}_{t}$ yielding a higher utility than ${\bm{x}}^{\prime}_{t}$. While this type of feedback is more consistent in practice, it creates quite a challenging problem compared to the logistic problem of Section˜4. The search space for action pairs ${\mathcal{X}}\times{\mathcal{X}}$ is significantly larger than $\mathcal{X}$}, and the observed preference feedback of $s(f({\bm{x}}_{t})-f({\bm{x}}^{\prime}_{t}))$ conveys only relative information between two actions. We start by presenting a solution to estimate $f$ and obtain valid confidence sets under preference feedback. Using these confidence sets we then design the MaxMinLCB algorithm which chooses action pairs that are not only favorable, i.e., yield high utility, but are also informative for improving the utility confidence sets.
@@ 5.1 Preference-based Confidence Sets
We consider the probabilistic model of $y_{t}$ as stated in Section˜3, and write the corresponding regularized negative loglikelihood loss as
@@
This loss may be optimized over different function classes and is commonly used for linear dueling bandits (e.g., saha2021optimal), and has been notably successful in reinforcement learning with human feedback (christiano2017deep). We proceed to show that the preference-based loss ${\mathcal{L}}_{k}^{\mathrm{D}}$ is equivalent to ${\mathcal{L}}_{k^{\mathrm{D}}}^{\mathrm{L}}$, the standard logistic loss (3) invoked with a specific kernel function $k^{\mathrm{D}}$. This will allow us to cast the problem of inference with preference feedback as a kernelized logistic regression problem. To this end, we define the dueling kernel as
@@
The proof is left to Section˜C.1. The assumption on eigenfunctions in Proposition˜4 is primarily made to simplify the equivalence class. In particular, the relative preference function $h$ can only capture the utility $f$ up to a bias, i.e., if a constant bias $b$ is added to all values of $f$, the corresponding $h$ function will not change. The value of $b$ may not be recovered by drawing queries from $h$, however, this will not cause issues in terms of identifying $\operatorname*{arg\,max}$ of $f$ through querying values of $h$. Therefore, we set $b=0$ by assuming that eigenfunctions of $k$ are zero-mean. This assumption automatically holds for all kernels that are translation or rotation invariant over symmetric domains, since their eigenfunctions are periodic $L_{2}({\mathcal{X}})$ basis functions, e.g., Matérn kernels and sinusoids.
Proposition˜4 allows us to re-write the preference-based loss function of (6) as a logistic-type loss
@@
Assume $f\in{\mathcal{H}}_{k}$ and $\left\lVert f\right\rVert_{k}\leq B$. Choose $0<\delta<1$ and set $\beta^{\mathrm{D}}_{t}(\delta)$ and $\sigma^{\mathrm{D}}_{t}$ as in equations (4) and (5), with $k^{D}$ used as the kernel function. Then,
@@
Corollary˜5 gives valid confidence sets for kernelized utility functions under preference feedback and immediately improves prior results on linear dueling bandits and kernelized dueling bandits with regression-type loss, to kernelized setting with logistic-type likelihood. To demonstrate this, in Section˜C.3 we present the kernelized extensions of MaxInP (saha2021optimal, Algorithm 3), and IDS (kirschner2021bias, Algorithm 4) and prove the corresponding regret guarantees (cf. Theorems 15 and 16). Corollary˜5 holds almost immediately by invoking Theorem˜2 with the dueling kernel $k^{\mathrm{D}}$ and applying Proposition˜4. A proof is provided in Section˜C.1 for completeness.
Comparison to Prior Work. A line of previous work assumes that both $f$ and the probability $s(f({\bm{x}}))$ are $B$-bounded members of ${\mathcal{H}}_{k}$. This allows them to directly estimate $s(f({\bm{x}}))$ via kernelized linear regression (xu2020zeroth; mehta2023kernelized; kirschner2021bias). The resulting confidence intervals are then around the least squares estimator, which does not align with the logistic estimator $f_{t}$}. This model does not encode the fact that $s(f({\bm{x}}))$ only takes values in $[0,1]$ and considers a sub-Gaussian distribution for $y_{t}$, instead of the Bernoulli formulation when calculating the likelihood. Therefore, the resulting algorithms require more samples to learn an accurate reward estimate. In a concurrent work, xu2024principled also consider the loss function of Equation˜6 and present likelihood-ratio confidence sets. The width of the sets at time $T$, scales with $\sqrt{T\log{\mathcal{N}}({\mathcal{H}}_{k};1/T})$ where the second term is the metric entropy of the $B$-bounded RKHS at resolution $1/T$, that is, the log-covering number of this function class, using balls of radius $1/T$. It is known that $\log{\mathcal{N}}({\mathcal{H}}_{k};1/T)\asymp\gamma_{T}$ as defined in Theorem˜2. This may be easily verified using wainwright2019high and (vakili2021information, Definition 1). Noting the definition of $\beta_{t}^{\mathrm{D}}$, we see that likelihood ratio sets of xu2024principled are wider than Corollary˜5. Consequently, the presented regret guarantee in this work is looser by a factor of $T^{1/4}$ compared to our bound in Theorem˜6.
@@
$\text{s.t. }{\bm{x}}^{\prime}({\bm{x}})$ $=\arg\min_{{\bm{x}}^{\prime}\in{\mathcal{M}}_{t}}s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))-\beta^{\mathrm{D}}_{t}\sigma^{\mathrm{D}}_{t}({\bm{x}},{\bm{x}}^{\prime})$
$\text{and }{\bm{x}}_{t}^{\prime}$ $={\bm{x}}^{\prime}({\bm{x}}_{t}).$
@@ 5.2 Action Selection Strategy
Let ${\bm{x}}^{\prime}({\bm{x}})=\arg\min_{{\bm{x}}^{\prime}\in{\mathcal{M}}_{t}}\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})$ denote a response function. We propose MaxMinLCB in Algorithm˜1 for the preference feedback bandit problem that selects ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ jointly as
@@
where the lower-confidence bound $\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})=s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))-\beta^{\mathrm{D}}_{t}\sigma^{\mathrm{D}}_{t}({\bm{x}},{\bm{x}}^{\prime})$ presents a pessimistic estimate of $h$ and ${\mathcal{M}}_{t}=\{{\bm{x}}\in{\mathcal{X}}|\,\forall{\bm{x}}^{\prime}\in{\mathcal{X}}:\,s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))+\beta^{\mathrm{D}}_{t}\sigma^{\mathrm{D}}_{t}({\bm{x}},{\bm{x}}^{\prime})\geq 0.5\}$ is the set of potentially optimal actions. The second action is chosen as ${\bm{x}}_{t}^{\prime}={\bm{x}}^{\prime}({\bm{x}}_{t})$. Equation˜7 forms a zero-sum Stackelberg (Leader–Follower) game where the actions ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ are chosen sequentially (stackelberg1952theory). First, the Leader selects ${\bm{x}}_{t}$, then the Follower selects ${\bm{x}}_{t}^{\prime}$ depending on the choice of ${\bm{x}}_{t}$. Importantly, due to the sequential nature of action selection, ${\bm{x}}_{t}$ is chosen by the Leader such that the Follower’s action selection function, ${\bm{x}}^{\prime}(\cdot)$, is accounted for in the selection of ${\bm{x}}_{t}$. Sequential optimization problems are known to be computationally NP-hard even for linear functions (jeroslow1985polynomial). However, due to their importance in practical applications, there are algorithms that can efficiently approximate a solution over large domains (sinha2017review; ghadimi2018approximation; dagreou2022framework; camacho2023metaheuristics).
@@
MaxMinLCB builds on a simple insight: if the utility $f$ is known, both the Leader and the Follower will choose $\bm{x}$$^{\star}$ yielding an objective value $0.5$ for both players, and zero dueling regret. Since MaxMinLCB has no access to $f$, it leverages the confidence sets of Corollary˜5 and uses a pessimistic approach by considering the LCB instead. There are two crucial properties of the Follower specific to this game. First, the Follower can not do worse than the Leader with respect to the $\mathrm{LCB}_{t}$. In any scenario, the Follower can match the Leader’s action which results in $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})=0.5$. Second, for sufficiently tight confidence sets, the Follower will not select sub-optimal actions. In this case, the Leader’s best action must be optimal as it anticipates the Follower’s response and Equation˜7 recovers the optimal actions. Therefore, the objective value of the game considered in Equation˜7 is always less than, or equal to the objective of the game with known utility function $f$, i.e., $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})\leq 0.5=f({\bm{x}}^{\star},{\bm{x}}^{\star})$ and the gap shrinks with the confidence sets. Overall, the Stackelberg game in Equation˜7 can be considered as a lower approximation of the game played with known utility function $f$.
@@
The primary challenge for MaxMinLCB is to sample action pairs that sufficiently shrink the confidence sets for the optimal actions without accumulating too much regret. MaxMinLCB balances this exploration-exploitation trade-off naturally with its game theoretic formulation. We view the selection of ${\bm{x}}_{t}$ to be exploitative by trying to maximize the unknown utility $f({\bm{x}}_{t})$ and minimizing regret. On the other hand, ${\bm{x}}_{t}^{\prime}$ is chosen to be the most competitive opponent to ${\bm{x}}_{t}$, i.e., testing whether the condition $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})\geq 0.5$ holds. Note that $\mathrm{LCB}_{t}$ is pessimistic concerning ${\bm{x}}_{t}$ making it robust against the uncertainty in the confidence set estimation. At the same time, $\mathrm{LCB}_{t}$ is an optimistic estimate for ${\bm{x}}_{t}^{\prime}$ encouraging exploration. In our main theoretical result, we prove that under the assumptions of Corollary˜5, MaxMinLCB achieves sublinear regret on the dueling bandit problem.
@@
Suppose the utility function $f$ lies in ${\mathcal{H}}_{k}$ with a norm bounded by $B$, and that kernel $k$ satisfies the assumption of Proposition˜4. Let $\delta\in(0,1]$ and choose the exploration coefficient $\beta^{\mathrm{D}}_{t}(\delta)$ as in Corollary˜5. Then MaxMinLCB satisfies the anytime dueling regret of
@@
The proof is left to Section˜C.2. The information gain $\gamma_{T}^{\mathrm{D}}$ in Theorem˜6 quantifies the structural complexity of the RKHS corresponding to $k^{\mathrm{D}}$ and its dependence on $T$ is fairly understood for kernels commonly used in applications of bandit optimization. As an example, for a Matérn kernel of smoothness $\nu$ defined over a $d$-dimensional domain, $\gamma_{T}=\tilde{\mathcal{O}}(T^{d/(2\nu+d)})$ (Remark 2, vakili2021information) and the corresponding regret bound grows sublinearly with $T$.
Restricting the optimization domain to ${\mathcal{M}}_{t}\subset{\mathcal{X}}$ is common in the literature (zoghi2014relative; saha2021optimal) despite being challenging in applications with large or continuous domains. We conjecture that MaxMinLCB would enjoy similar regret guarantees without restricting the selection domain to ${\mathcal{M}}_{t}$ as done in Equation˜7. This claim is supported by our experiments in Section˜6.2 which are carried out without such restriction on the optimization domain.
@@
Figure 1: Regret of learning the Ackley function with logistic and preference feedback. (a) Same UCB algorithms, each using a different confidence set. LGP-UCB performs best, showcasing the power of Theorem˜2. (b): Algorithms with different acquisition functions, all using our confidence sets. MaxMinLCB is more sample-efficient.
@@ 6 Experiments
Our experiments are on finding the maxima of test functions commonly used in (non-convex) optimization literature (jamil2013literature), given only preference feedback. These functions cover challenging optimization landscapes including several local optima, plateaus, and valleys, allowing us to test the versatility of MaxMinLCB. We use the Ackley function for illustration in the main text and provide the regret plots for the remainder of the functions in Appendix˜E. For all experiments, we set the horizon $T=2000$ and evaluate all algorithms on a uniform mesh over the input domain of size $100$. Additionally, we conducted experiments on the Yelp restaurant review dataset to demonstrate the applicability of MaxMinLCB on real-world data and its scaling to larger domains. All experiments are run across $20$ random seeds and reported values are averaged over the seeds, together with standard error. The environments and algorithms are implemented [^2] The code is made available at github.com/lasgroup/MaxMinLCB. end-to-end in JAX (jax2018github).